OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike? I thought you would start to solve it by putting the blue bike in terms of red so.. b=14-r and b+r=300 so then 14-r+r=300 thats where i get stuck.. help!! A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike? I thought you would start to solve it by putting the blue bike in terms of red so.. b=14-r and b+r=300 so then 14-r+r=300 thats where i get stuck.. help!! @Mathematics

OpenStudy (anonymous):

b is not equal to 14-r b is r-14

OpenStudy (lilg132):

b = r - 14 and b + r = 300

OpenStudy (lilg132):

(r - 14) + r = 300

OpenStudy (anonymous):

ok i understand now thanks!

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

red bike = $157 blue bike = $ 143

OpenStudy (anonymous):

150+16 = 166 not 154

OpenStudy (anonymous):

sorry not 164

OpenStudy (anonymous):

i meant to type 14 not 16

myininaya (myininaya):

R=B+14 R+B=300

myininaya (myininaya):

\[(B+14)+B=300\]

myininaya (myininaya):

\[2B=300-14\]

myininaya (myininaya):

\[B=150-7=143\]

OpenStudy (anonymous):

red bike = $157 blue bike = $ 143

myininaya (myininaya):

R=B+14=143+14=157 yes i agree

OpenStudy (lilg132):

300 - 14 = 286 286 / 2 = 143 (blue bike) 300 - 143 = 157 (red bike)

OpenStudy (phi):

When you see a statement like A blue bike is $14 less than a red bike. interpret the "X less than" as -X + B is 14 less than R B= -14 + R which you can re-write as B= R-14

myininaya (myininaya):

or you can read it as the red bike is $14 more than the blue bike :)

myininaya (myininaya):

R=B+14